OpenStudy (alana242):

The difference between two numbers is 9. Find the numbers. If their sum is 46

OpenStudy (studygurl14):

x and y are the two numbers y - x = 9 x + y = 46 Solve the system of equations

OpenStudy (alana242):

Why would you use y though . I'm so confused. I know the answer is 5 but idk how to get there

OpenStudy (studygurl14):

It doesn't matter what variables you use.

OpenStudy (mathmale):

It's just about essential that you choose variables, such as x and y, to represent your unknowns. y-x=9 states that "the difference between the unknown numbers x and y is 9."

OpenStudy (studygurl14):

^Like @mathmale stated. It is just important that you use them, not which letters you pick

OpenStudy (mathmale):

Let's take: y - x = 9 x + y = 46 and rewrite the 2nd equation: y - x = 9 y + x = 46 --------- what would happen if you were now to add the 2 equations together? Can you find y?

OpenStudy (studygurl14):

@alana242